# On the PM2.5 – Mortality Relationship: A Bayesian Model for Spatio-Temporal Confounding

### v2

@@ 3 The Proposed Model
It is assumed that the outcome, $Y_{it}$, follows a Poisson distribution with mean $\mu_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$. The linear predictor, $\vartheta_{it}=\ln(\mu_{it})$, is modeled as a function of exposure, $X_{it}$, measured confounders, $\mathbf{M}_{it}=({M}_{1it},\dots,{M}_{{p_{m}}it})^{\prime}$, and an unmeasured confounder, $Z_{it}$, which is a smooth function of space and time. Assuming that ${\mathbf{M}}_{it}$ and $Z_{it}$ are spatially and temporally smoother than the exposure, the model can be specified as follows:
@@
$\vartheta_{it}$ $={\color[rgb]{0,0,0}\ln(\mu_{it})}=o_{it}+\beta_{0}+f(X_{it};\bm{\gamma}_{f})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+u_{it}\,,$ (2)
where $o_{it}=\log(E_{it})$ is the offset, $\beta_{0}$ is an intercept, $f(\cdot;\cdot)$ is an unknown smooth function parameterized by the vector $\bm{\gamma}_{f}$, $\bm{\alpha}$ is a ${p_{m}}$-dimensional vector of coefficients, and $u_{it}$ is a spatially and temporally uncorrelated Gaussian random variable, assumed to have zero mean and constant variance, i.e. $u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})$.
If all confounders were measured and $f(\cdot;\cdot)$ was modeled using basis functions, the model specification in Equations (1)–(2) would belong to the family of generalized additive models (GAMs; see Wood, 2017). Alternatively, considering that $f(\cdot;\cdot)$ can be adequately approximated by using the first-order Taylor’s expansion, we have:
@@
where $\widehat{X}_{it}$ denotes the large-scale (or overall) spatio-temporal trend, $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the first derivative of $f(\cdot;\bm{\gamma}_{f})$ with respect to its first argument, evaluated at $\widehat{X}_{it}$, and $\varpi_{it}$ denotes the Taylor’s remainder. To calculate the spatio-temporal trend, $X_{it}$ can be regressed on a smooth function of space and time (see also Janes et al., 2007):
${\color[rgb]{0,0,0}X_{it}=h({\bf s}_{i},t;df)+\nu_{it}\,,}$ (4)
where ${\bf s}_{i}$ denotes the spatial coordinates of the centroid of areal unit $i$, $df$ are the degrees of freedom of the smooth term, and $\nu_{it}$ is an uncorrelated Gaussian random variable with zero mean and constant variance. The overall trend is then defined as the fitted values obtained from the model (4), i.e. $\widehat{X}_{it}=\widehat{h}({\bf s}_{i},t;df)$.
Equation (3) provides a decomposition of the exposure effect into two components describing large and small (or local) scales of variation. These components are functions of $\widehat{X}_{it}$ and $X_{it}-\widehat{X}_{it}$, respectively. They are orthogonal because $X_{it}-\widehat{X}_{it}$ is the residual term from the regression in Equation (4). Henceforth, the decomposition above should help to disentangle the effect of interest from that of unmeasured confounding factors. The rationale for using detrended exposure, as discussed in the econometric and spatio-temporal literature, is to remove long-term and large-scale spatial variation, preserving only short-term and local fluctuations. This helps reduce the risk of spurious associations driven by underlying trends (Wooldridge, 2020). From an epidemiological point of view, consider that $\widehat{X}_{it}$ captures trend and seasonal components of the exposure, therefore it is expected to be highly correlated with $Z_{it}$, given that the latter is smooth, while the correlation between $X_{it}-\widehat{X}_{it}$ and $Z_{it}$ should be close to zero. Therefore, the unmeasured confounder is more likely to affect the association between outcome and exposure at large scales (Greven et al., 2011, Janes et al., 2007), so the first term in the decomposition of $f(X_{it};\bm{\gamma}_{f})$ in Equation (3) might be biased, but it is less likely that the second term, $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})$, is affected by confounding bias.
By decomposing $f(X_{it};\bm{\gamma}_{f})$ in Equation (2), we thus have:
$\vartheta_{it}=o_{it}+\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+\tilde{u}_{it}\,.$ (5)
@@
For each $i$, the unknown function evaluated at the overall trend of the exposure, $f(\widehat{X}_{it};\bm{\gamma}_{f})$, and its first derivative in Equation (5) can be seen as functions of time, hence time-varying coefficients can adequately represent them (see also West and Harrison, 1997, Section 3.1). Therefore, we propose to replace the linear predictor in Equation (2) with:
@@
where $\beta_{1it}=f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the spacetime-varying coefficient representing the local association between exposure and outcome (in the sense of Janes et al. (2007)), while $\beta_{0it}$ is a spacetime-varying intercept,
$\beta_{0it}=\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+Z_{it}\,,$
that should account for both the large-scale exposure effect and the unmeasured confounder.
We now provide two examples of the function $f(\cdot;\cdot)$ and of the resulting varying coefficients in Equation (6). These show that our model generalizes other methods in the literature.
@@
If $f(\cdot;\bm{\gamma}_{f})$ is a constant function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}$, then $\beta_{1it}=0$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+Z_{it}$.
@@
If $f(\cdot;\bm{\gamma}_{f})$ is a linear function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}$, then $\beta_{1it}=\gamma_{f1}$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+\gamma_{f1}\widehat{X}_{it}+Z_{it}$. The model proposed by Greven et al. (2011) can be cast in this case, with $\gamma_{f0}=0$. The authors, however, assume further that the intercepts do not depend on time.
@@ 3.1 Specification of the Spacetime-Varying Coefficients
Although it is possible to model $\beta_{0it}$ and $\beta_{1it}$ directly, we decompose these spacetime-varying coefficients as follows to enhance prior elicitation and interpretability (see also Bakar et al., 2015, Cai et al., 2013, Paez et al., 2008):
$\beta_{kit}=\overline{\delta}_{k}+\delta_{ki}+\delta_{kt}+\delta_{kit}^{*}\,,\quad k=0,1\,,$
where $\overline{\delta}_{k}$ is the global effect of the $k$th covariate over the whole study period, $\delta_{ki}$ is the area-specific effect of the $k$th covariate, $\delta_{kt}$ is the time-specific effect of the $k$th covariate, and $\delta_{kit}^{*}$ is the spatio-temporal interaction effect of the $k$th covariate. For each $i$, $\delta_{ki}$ can be interpreted as the average (across time) spatial deviation from the global effect. In particular, $\delta_{1i}$ captures time-independent spatial heterogeneity in the exposure–outcome association at the local scale of variation (recall that the regressor is $X_{it}-\widehat{X}_{it}$). It can inform about where populations are most vulnerable to the effects of air pollution.
For each time instant, $\delta_{kt}$ can be interpreted as the average (across space) temporal changes from the global effect. Hence, $\delta_{1t}$ represents the temporal evolution of the association of interest, and highlights when peaks can occur. The interaction effect, $\delta_{kit}^{*}$, captures the area- and time-specific deviations from the global effect that are left after accounting for the temporal and spatial effects. The simulation scenarios in Section 6 were built based on possible values of $\beta_{1it}$ that are very small in real applications. Hence, given the tiny contribution of the term $\beta_{1it}(X_{it}-\widehat{X}_{it})$ to the augmented-data likelihood (see Section 4.2), it is difficult to recover the interaction effect of the exposure, even in the absence of confounding. Therefore, we assume that $\delta_{1it}^{*}$ is zero. This assumption is in line with previous studies (see, e.g., Bakar et al., 2015, Paez et al., 2008).
In SDGLMC, we implement a DGLM with random walk (RW) processes to model the temporal effects, $\delta_{0t}$ and $\delta_{1t}$, and the interaction effect, $\delta_{0it}^{*}$. We propose the following state-space formulation for the latent process and the regression coefficients, for $t=1,\dots,{T}$:
@@
where the symbol $\odot$ denotes the Hadamard product, $\bm{\vartheta}_{t}=(\vartheta_{1t},\dots,\vartheta_{{N}t})^{\prime}$, $\mathbf{o}_{t}=(o_{1t},\dots,o_{{N}t})^{\prime}$, $\tilde{\bm{\delta}}_{k}=(\delta_{k1},\dots,\delta_{k{N}})^{\prime}$ are the spatial-specific effects for $k=0,1$, $\tilde{{\bf u}}_{t}=(\tilde{u}_{1t},\dots,\tilde{u}_{{N}t})^{\prime}\sim\mathcal{N}({\bf 0},\tilde{\tau}_{u}^{2}{\bf I}_{N})$, $\bm{\delta}_{0t}^{*}=(\delta_{01t}^{*},\dots,\delta_{0{N}t}^{*})^{\prime}$, and the observation matrices are as follows:
@@
Note that (9) is one of five alternatives that will be discussed further in Section 7. Similarly to $\mathbf{w}_{0t}^{*}$, to capture the spatial correlation among the area-specific effects, $\tilde{\bm{\delta}}_{0}$ and $\tilde{\bm{\delta}}_{1}$ are modeled as intrinsic conditional autoregressive (ICAR; see, e.g., Banerjee et al., 2014) processes, denoted as $ICAR(\sigma_{\delta k}^{2})$, with density proportional to
@@
where $\mathbf{W}$ is the spatial adjacency matrix with elements $(\mathbf{W})_{ij}=1$ if areal units $i$ and $j$ are neighbors, and 0 otherwise, $\mathbf{D}$ is a diagonal matrix with the sum of the rows of $\mathbf{W}$, and $\sigma_{\delta k}^{2}$ is a variance parameter. Such structure has been previously considered by Franco-Villoria et al. (2019, 2022) to model spatially varying coefficients.
@@ 4 Bayesian Inference and Computations in SDGLMC
The first stages of the Bayesian hierarchical model are defined by Equations (1) and (7)–(10). To complete it, a prior distribution must be assigned on the parameter set, $\bm{\Theta}=\{\overline{\delta}_{0},\delta_{00},\bm{\delta}_{00}^{*},\overline{\delta}_{1},\delta_{10},\bm{\alpha},\sigma_{w0}^{2},\sigma_{\delta 0}^{2},\sigma_{w0}^{*2},\sigma_{w1}^{2},\sigma_{\delta 1}^{2},\tilde{\tau}_{u}^{2}\}$.
@@ 4.1 Prior Specification
Independent prior distributions are imposed on each parameter in $\bm{\Theta}$. As for the time-varying coefficients, distributional assumptions are required on the initial states (i.e., when $t=0$), and on the variances of innovation errors. We impose that $\delta_{k0}\sim\mathcal{N}(0,V_{\delta k})$, $k=0,1$, where $V_{\delta 0}=4$ and $V_{\delta 1}=4\times 10^{-6}$. Also, $\bm{\delta}_{00}^{*}\sim\mathcal{N}(\mathbf{0},\mathbf{V}_{\delta}^{*})$, with $\mathbf{V}_{\delta}^{*}=10\mathbf{I}_{N}$.
Inverse-gamma priors are imposed on $\sigma_{w0}^{*2}\sim\mathcal{IG}(a_{w0}^{*},b_{w0}^{*})$, and similarly on $\sigma_{wk}^{2}\sim\mathcal{IG}(a_{wk},b_{wk})$ and $\sigma_{\delta k}^{2}\sim\mathcal{IG}(a_{\delta k},b_{\delta k})$, for $k=0,1$. For $k=0$, all the hyperparameters are assumed to be $0.01$. For $k=1$, choosing the hyperparameters requires more care. Recall that capturing gradual changes is a desired feature of the proposed model, and the variances $\sigma_{wk}^{2}$ and $\sigma_{\delta 1}^{2}$ play a crucial role to this end. For instance, if their prior means are “large”, the variations in the temporal and spatial effects, respectively, become unlikely pronounced, whereas if their prior means are “small”, the “true” variation is captured only in part. The definition of “large” and “small” mainly depends on the sample size and on how much the prior dominates the likelihood.
Therefore, several values for the hyperparameters in both inverse-gamma distributions are considered to evaluate the sensitivity of the posterior distribution to the prior. In the case of both simulated and real data (see Sections 6 and 7, respectively), we find that $a_{w1}=10$, $b_{w1}=10^{-3}$, and $a_{\delta 1}=1,b_{\delta 1}=0.01$ provide a good balance between the prior and the likelihood.
Finally, the global effects, $\overline{\delta}_{0}$ and $\overline{\delta}_{1}$, and the coefficients of the measured confounders, $\bm{\alpha}$, are assumed independently and normally distributed with mean zero and variance $10^{6}$. Following Wakefield (2007), it is assumed that the residual relative risks, $\tilde{u}_{it}$, follow a log-t distribution with two degrees of freedom, and with $95\%$ of prior mass within the interval $(0.98,1.02)$, hence we obtain the prior $\tilde{\tau}_{u}^{2}\sim\mathcal{IG}(1,2.1\times 10^{-5})$.

### v3

@@ 3 The Proposed Model
It is assumed that the outcome, $Y_{it}$, follows a Poisson distribution with mean $\mu_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$. The linear predictor, $\vartheta_{it}=\ln(\mu_{it})$, is modeled as a function of exposure $X_{it}$, measured confounders $\mathbf{M}_{it}=({M}_{1it},\dots,{M}_{{p_{m}}it})^{\prime}$, and an unmeasured confounder $Z_{it}$, which is smooth in space and time. Assuming that ${\mathbf{M}}_{it}$ and $Z_{it}$ are spatially and temporally smoother than the exposure, the model is specified as follows:
@@
$\vartheta_{it}$ $=\ln(\mu_{it})=o_{it}+\beta_{0}+f(X_{it};\bm{\gamma}_{f})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+u_{it}\,$, (2)
where $o_{it}=\log(E_{it})$ is the offset, $\beta_{0}$ is an intercept, $f(\cdot;\cdot)$ is an unknown smooth function parameterized by the vector $\bm{\gamma}_{f}$, $\bm{\alpha}$ is a $p_{m}$}-dimensional vector of coefficients, and $u_{it}$ is a spatially and temporally uncorrelated Gaussian error term with zero mean and constant variance, $u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})$.
If all confounders were measured and $f(\cdot;\cdot)$ was modeled using basis functions, the model specification in Equations (1) and (2) would belong to the class of generalized additive models (GAMs; see Wood, 2017). Instead, we consider that $f(\cdot;\cdot)$ can be adequately approximated by the first-order Taylor’s expansion:
@@
where $\widehat{X}_{it}$ denotes the large-scale (or overall) spatio-temporal trend, $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the first derivative of $f(\cdot;\bm{\gamma}_{f})$ with respect to its first argument evaluated at $\widehat{X}_{it}$, and $\varpi_{it}$ denotes the Taylor’s remainder. To compute the spatio-temporal trend, $X_{it}$ can be regressed on a smooth function of space and time (see also Janes et al., 2007):
$X_{it}=h({\bf s}_{i},t;df)+\nu_{it}\,$, (4)
where ${\bf s}_{i}$ denotes the spatial coordinates of the centroid of areal unit $i$, dfdf are the degrees of freedom of the smooth term, and $\nu_{it}$ is an uncorrelated Gaussian random variable with zero mean and constant variance. The overall trend is then defined as the fitted values obtained from the model (Equation (4)), i.e., $\widehat{X}_{it}=\widehat{h}({\bf s}_{i},t;df)$.
Equation (3) thus decomposes the exposure effect into two orthogonal components: one at large scale of variation - $f(\widehat{X}_{it};\bm{\gamma}_{f})$ - and one at a small (or local) scale - $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})$. This decomposition helps disentangle the effect of interest from that of unmeasured confounders. As discussed in the econometric and spatio-temporal literatures, the use of detrended exposure helps remove large-scale spatio-temporal variation, thereby reducing the risk of spurious associations driven by underlying trends (Wooldridge, 2020). At the same time, it preserves small-scale spatio-temporal fluctuations. From an epidemiological point of view, $\widehat{X}_{it}$ captures trend and seasonal components of the exposure and, thus, is expected to be highly correlated with $Z_{it}$ (which is assumed to be smooth), whereas the correlation between $X_{it}-\widehat{X}_{it}$ and $Z_{it}$ should be close to zero. Therefore, the unmeasured confounder conceivably affects the association between outcome and exposure at large scales (Greven et al., 2011, Janes et al., 2007), so the first term in Equation (3) can potentially be biased, but it is less likely that the second term $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})$ is affected by confounding bias.
Substituting the decomposition into Equation (2), we obtain:
$\vartheta_{it}=o_{it}+\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+\tilde{u}_{it}\,$,
@@
For each $i$, the unknown function $f(\widehat{X}_{it};\bm{\gamma}_{f})$ and its first derivative can be seen as functions of space and time, and spacetime-varying coefficients can adequately represent them (see also West and Harrison, 1997, Section 3.1). Therefore, we propose the following reparameterization:
@@
where $\beta_{1it}=f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the spacetime-varying coefficient representing the local association between exposure and outcome. It denotes the association between day-to-day variation in area-specific deviations in PM2.5 concentrations from the large-scale spatio-temporal trend, and the corresponding day-to-day variation in area-specific deviations in mortality rates from the large-scale trend. The parameter $\beta_{0it}$ is a spacetime-varying intercept that accounts for both the large-scale exposure effect and the unmeasured confounder:
$\beta_{0it}=\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+Z_{it}\,.$
The following examples illustrate that under certain function forms of $f(\cdot;\cdot)$, our model extends other methods proposed in the literature.
@@
If $f(\cdot;\bm{\gamma}_{f})$ is a constant function, i.e., $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}$, then $\beta_{1it}=0$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+Z_{it}$.
@@
If $f(\cdot;\bm{\gamma}_{f})$ is a linear function, i.e., $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}$, then $\beta_{1it}=\gamma_{f1}$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+\gamma_{f1}\widehat{X}_{it}+Z_{it}$. The model proposed by Greven et al. (2011) can be cast in this case, with $\gamma_{f0}=0$. The authors, however, assume further that the intercepts do not depend on time.
@@ 3.1 Specification of the Spacetime-Varying Coefficients
Although it is possible to model $\beta_{0it}$ and $\beta_{1it}$ directly, we decompose them to enhance prior elicitation and interpretability (see also Bakar et al., 2015, Cai et al., 2013, Paez et al., 2008):
$\beta_{kit}=\overline{\delta}_{k}+\tilde{\delta}_{ki}+\delta_{kt}+\delta_{kit}^{*}\,,\quad k=0,1\,$,
where $\overline{\delta}_{k}$ is the baseline effect of the $k$th covariate, which is constant across space and time; $\tilde{\delta}_{ki}$ is the area-specific effect of the $k$th covariate; $\delta_{kt}$ is the time-specific effect of the $k$th covariate; and $\delta_{kit}^{*}$ is the spatio-temporal interaction effect of the $k$th covariate. For each $i$, $\tilde{\delta}_{ki}$ represents the average (across time) spatial deviation from the baseline effect. In particular, $\tilde{\delta}_{1i}$ captures time-independent spatial heterogeneity in the exposure–outcome association at the small scale of variation (recall that the regressor is $X_{it}-\widehat{X}_{it}$). Thus, the parameter $\tilde{\delta}_{1i}$ can inform about where populations are most vulnerable to air pollution.
For each time $t$, $\delta_{kt}$ can be interpreted as the average (across space) temporal deviations from the baseline effect. Thus, $\delta_{1t}$ represents the temporal evolution of the association of interest, and highlights when peaks can occur.
The interaction effect, $\delta_{kit}^{*}$, captures the area- and time-specific deviations that are left after accounting for the temporal and spatial effects. The simulation scenarios in Section 6 are built based on plausible values of $\beta_{1it}$, which are typically very small in empirical applications. Thus, given the tiny contribution of the term $\beta_{1it}(X_{it}-\widehat{X}_{it})$ to the augmented-data likelihood (see Section 4.2), it is difficult to recover the interaction effect of the exposure, even in the absence of confounding. Therefore, we posit that $\delta_{1it}^{*}$ equals zero, an assumption consistent with prior research (e.g., Bakar et al., 2015, Paez et al., 2008).
In SDGLMC, we implement a DGLM with random walk (RW) processes to model the temporal effects (namely, $\delta_{0t}$ and $\delta_{1t}$) and the interaction effect, $\delta_{0it}^{*}$. We propose the following state-space formulation for the latent process and the regression coefficients, for $t=1,\dots,{T}$:
@@
where the symbol $\odot$ denotes the Hadamard product, $\bm{\vartheta}_{t}=(\vartheta_{1t},\dots,\vartheta_{{N}t})^{\prime}$, $\mathbf{o}_{t}=(o_{1t},\dots,o_{{N}t})^{\prime}$, $\tilde{\bm{\delta}}_{k}=(\tilde{\delta}_{k1},\dots,\tilde{\delta}_{k{N}})^{\prime}$ are the spatial-specific effects for $k=0,1$, $\tilde{{\bf u}}_{t}=(\tilde{u}_{1t},\dots,\tilde{u}_{{N}t})^{\prime}\sim\mathcal{N}({\bf 0},\tilde{\tau}_{u}^{2}{\bf I}_{N})$, $\bm{\delta}_{0t}^{*}=(\delta_{01t}^{*},\dots,\delta_{0{N}t}^{*})^{\prime}$, and the observation matrices are as follows:
@@
Note that Equation (8) is one of five alternatives that will be discussed further in Section 7. To capture the underlying spatial structure, $\tilde{\bm{\delta}}_{0}$ and $\tilde{\bm{\delta}}_{1}$ are modeled as intrinsic conditional autoregressive (ICAR; e.g., Banerjee et al., 2014) processes, denoted as $ICAR(\sigma_{\delta k}^{2})$, with probability density function proportional to
@@
where $\mathbf{W}$ is the spatial adjacency matrix with elements $(\mathbf{W})_{ij}=1$ if areal units $i$ and $j$ are neighbors, and 0 otherwise; $\mathbf{D}$ is a diagonal matrix constructed from the sum of the rows of $\mathbf{W}$; and $\sigma_{\delta k}^{2}$ is a variance parameter. Such structure has been previously considered by Franco-Villoria et al. (2019, 2022) to model spatially varying coefficients.
@@ 4 Bayesian Inference and Computations in SDGLMC
The first stages of the Bayesian hierarchical model are defined by Equations (1) and (6)–(9). To complete these stages, we assign a prior distribution on the parameter set, $\bm{\Theta}=\{\overline{\delta}_{0},\delta_{00},\bm{\delta}_{00}^{*},\overline{\delta}_{1},\delta_{10},\bm{\alpha},\sigma_{w0}^{2},\sigma_{\delta 0}^{2},\sigma_{w0}^{*2},\sigma_{w1}^{2},\sigma_{\delta 1}^{2},\tilde{\tau}_{u}^{2}\}$.
@@ 4.1 Prior Specification
Independent prior distributions are imposed on each component of $\bm{\Theta}$. It is assumed that the initial states of the time-varying coefficients follow a Gaussian distribution, namely $\delta_{k0}\sim\mathcal{N}(0,V_{\delta k})$, $k=0,1$, where $V_{\delta 0}=4$ and $V_{\delta 1}=4\times 10^{-6}$. Also, $\bm{\delta}_{00}^{*}\sim\mathcal{N}(\mathbf{0},\mathbf{V}_{\delta}^{*})$, with $\mathbf{V}_{\delta}^{*}=10\mathbf{I}_{N}$.
Inverse-gamma priors are imposed on the variances of the innovation errors, namely $\sigma_{w0}^{*2}\sim\mathcal{IG}(a_{w0}^{*},b_{w0}^{*})$, $\sigma_{wk}^{2}\sim\mathcal{IG}(a_{wk},b_{wk})$ and $\sigma_{\delta k}^{2}\sim\mathcal{IG}(a_{\delta k},b_{\delta k})$, for $k=0,1$. For $k=0$, all the hyperparameters are assumed to be $0.01$. For $k=1$, care must be taken when specifying the hyperparameters because capturing gradual changes is a desired feature of the proposed model. If non-informative prior distributions are imposed on the variances, $\sigma_{w1}^{2}$ and $\sigma_{\delta 1}^{2}$, the variations in the corresponding temporal and spatial effects become unlikely pronounced. Therefore, a prior sensitivity analysis was conducted, where the hyperparameters $a_{w1}$, $b_{w1}$, $a_{\delta 1}$ and $b_{\delta 1}$ were varied to obtain a posterior distribution concentrated around relatively small values. The interpretation of “small” in this context is contingent upon the sample size and the extent to which the prior distribution influences the posterior inference relative to the likelihood. In the case of both simulated and real data (see Sections 6 and 7, respectively), we find that assuming $a_{w1}=10$, $b_{w1}=0.001$, $a_{\delta 1}=1$, and $b_{\delta 1}=0.01$ prevents overfitting as a priori $\Pr(\sigma_{w1}^{2}\leq 2.4\times 10^{-4})=0.99$ and $\Pr(\sigma_{\delta 1}^{2}\leq 0.995)=0.99$.
The baseline effects $\overline{\delta}_{0}$ and $\overline{\delta}_{1}$ and the coefficients of the measured confounders $\bm{\alpha}$ are assumed to be independently and normally distributed with mean zero and variance $10^{6}$. Finally, we assume that $\tilde{\tau}_{u}^{2}\sim\mathcal{IG}(1,2.1\times 10^{-5})$. These hyperparameters are chosen to reflect the belief that the residual relative risks $\exp(\tilde{u}_{it})$ follow a log-t distribution with two degrees of freedom, and with $95\%$ of prior mass within the interval $(0.98,1.02)$, as suggested by Wakefield (2007).